OpenStudy (anonymous):

A stone is dropped into a lake, creating a circular ripple that travels outward at a speed of 40 cm/s. Find the rate at which the area within the circle is increasing at 2s. 3s. & 7s.

OpenStudy (precal):

Isn't this a related rates problem involving calculus......?

OpenStudy (anonymous):

\[\frac{dA}{dt}=2\pi r\frac{dr}{dt}\] r at 2 seconds = 80cm \[\frac{dA}{dt}=2\pi (80cm)(40cm/s)\] do the same for r at 3seconds & 7 seconds